OpenStudy (anonymous):

Consider dy/dx = 1/2x + y -1 a) Find the values of the constants m and b, for which y = mx+ b is a solution to the differential equation.

OpenStudy (anonymous):

I got the same thing but I can't find out the value of b or m?

OpenStudy (anonymous):

There's another part of the question before this one. But I don't know if it's related to solving this one. I'll put in out just in case. Let y = f(x) be a particular solution to the differential equation with the initial condition f(0) = 1.

OpenStudy (anonymous):

let's plug in y(0) m=b-1

OpenStudy (anonymous):

Where did the x go?

OpenStudy (anonymous):

let's plug in y(1) m= 1/2 +(m+b) -1 1/2 + m+ b -1 two equation two unkown

OpenStudy (anonymous):

I plug in 0 for x in first one, 1 for x in second one

OpenStudy (anonymous):

\[m\to -\frac{1}{2},b\to \frac{1}{2}\]

OpenStudy (anonymous):

y= -1/2 x + 1/2 should work

OpenStudy (anonymous):

Thanks, I get it. But, if I face something on this level on my midterm, I think I'm still going to get roasted no matter how much I study.

OpenStudy (anonymous):

what class is that from?

OpenStudy (anonymous):

calculus

OpenStudy (anonymous):

oh, I've taken differentials, even then It took me some time as you can see

OpenStudy (anonymous):

There's a class called differentials? Wow. I'm just in plain calculus ab/bc. There's always some little holes I drop into when I do math.